# Virchow2: Scaling Self-Supervised Mixed Magnification Models in Pathology

### v2

@@ Abstract
Foundation models are rapidly being developed for computational pathology applications. However, it remains an open question which factors are most important for downstream performance with data scale and diversity, model size, and training algorithm all playing a role. In this work, we propose algorithmic modifications, tailored for pathology, and we present the result of scaling both data and model size, surpassing previous studies in both dimensions. We introduce two new models: Virchow2, a 632 million parameter vision transformer, and Virchow2G, a 1.9 billion parameter vision transformer, each trained with 3.1 million histopathology whole slide images, with diverse tissues, originating institutions, and stains. We achieve state of the art performance on 12 tile-level tasks, as compared to the top performing competing models. Our results suggest that data diversity and domain-specific methods can outperform models that only scale in the number of parameters, but, on average, performance benefits from the combination of domain-specific methods, data scale, and model scale.
@@ 1 Introduction
Figure 1: Virchow2 and Virchow2G adopt a pathology-adapted training method, scale up model and data size, and increase data diversity as compared to Virchow. Gains in the average weighted F1 score across 8 tile-level benchmarks using public data (see Tab. 3) are shown with respect to two scaling axes: model size (left), in terms of the number of parameters, and dataset size (right), in terms of the number of whole slide images (WSIs). Virchow2 improves over Virchow with pathology-specific modifications to the training method while scaling up the dataset from 1.5M to 3.1M WSIs, with increased diversity. Virchow2G then scales the model size from 632M to 1.9B parameters (ViT-H to ViT-G). Benchmark performance across foundation models in pathology appears to scale with model and dataset size.
@@ 3.1 Domain-specific augmentations
Figure 2: Highlighted modifications of Virchow tailored for pathology. a. Virchow2 and Virchow2G are trained with mixed magnification samples; slides are tiled at 40x, 20x, 10x, or 5x. b. Illustration of the difference between standard crop-and-resize augmentation (top) and the proposed extended-context translation (ECT) augmentation (bottom). ECT is expected to be more suitable for pathology images as it avoids distorting the cell morphology. We tuned the ECT approach to retain the average overlap between crops (shaded, teal). c. A diversity regularization objective seeks to push apart tile embeddings on a hypersphere (top). For Virchow2 and Virchow2G, we replace the unstable KoLeo regularizer used by Virchow with a kernel density estimator (KDE) that does not produce exploding gradients for similar tiles (bottom), which is particularly important for pathology whole slide images as they contain many highly correlated tiles.
@@ 5.5 Ablation results
Tab. 1 presents the results of the ablation analysis. In a majority of tasks (5/7), using all three domain-inspired changes, results in the best performance as well as the best average performance on in-distribution tasks. Using only ECT and KDE regularization has the best performance of 3/7 tasks and the highest average performance on out-of-distribution tasks, although the gap with all three is only 0.002. Interestingly, we observe that using only ECT leads to a small drop in performance while using only KDE regularization leads to a small improvement in performance on out-of-distribution tasks. Combining the two leads to a pronounced improvement, as noted above. All changes improved in-distribution tasks.
Geo.
Aug.
Phot.
Aug.
Reg.
PanMSK
20$\times$
PanMSK
10$\times$
PanMSK
5$\times$
PCam CRC MHIST MIDOG
ID
Avg.
OOD
Avg.
0.860 0.888 0.890 0.833 0.940 0.793 0.628 0.879 0.798
0.866 0.895 0.899 0.848 0.933 0.766 0.635 0.886 0.796
0.880 0.901 0.900 0.841 0.949 0.777 0.664 0.894 0.808
0.894 0.920 0.922 0.867 0.953 0.804 0.665 0.912 0.822
0.899 0.924 0.933 0.867 0.958 0.792 0.663 0.919 0.820
Table 1: Weighted F1 results for seven evaluation tasks and five model settings with a ViT-B/16. For each change, a checkmark indicates a deviation from the standard DINOv2 training protocol. In-distribution evaluation tasks include PanMSK XX, while out-of-distribution tasks include PCam, CRC, MHIST, and MIDOG. Boldface indicates the best performance and an underline indicates the second best performance. In a majority of tasks and on average, ECT, KDE, and no solarization regularization performs best.
@@ 6 Results of large-scale training
We improve on the training recipe of Virchow with pathology specific tuning and scale up the training first along the data axis (Virchow2) and then along the model axis (Virchow2G). We evaluate these models against other foundation models in pathology on tile-level benchmark tasks.
@@ 6.3 Embedding generation
For a 224 × 224 input tile image, a Virchow embedding is defined as the concatenation of the class token and the mean across all 256 of the patch tokens. This produces an embedding size of 2,560 (1,280 $\times$ 2) for Virchow, or 3,584 (1,792 $\times$ 2) for Virchow2 and Virchow2G. The class token is used for H-optimus-0 [45], Prov-GigaPath [58], UNI [13], and Phikon [25]. For CTransPath [54], the mean of all tokens is used as there is no class token.
@@ 6.4 Tile benchmarks
HEST-Benchmark is a public dataset of 74 spatial transcriptomics profiles from 48 patients grouped into 10 tasks based on organ [31]. The aim of the task is to predict the expression levels of the 50 most variable genes from 112 ×$\times 112\mu$m H&E stained image patches, each centered on a spatial transcriptomics spot. Here we present the results of a random forest classifier following the protocol described in Jaume et al. [31], as opposed to the linear protocol described in Sec. 5.3.
@@ 6.5 Scaling results
While we use the concatenation of the class [CLS] token with the mean of the patch tokens for all Virchow models, experiments are also performed using only the class token, with models denoted as “Virchow [CLS]”, “Virchow2 [CLS]”, and “Virchow2G [CLS]”. While Virchow2 [CLS] performs similarly to Virchow2, Virchow and Virchow2G benefit from the concatenated embedding. Using Virchow2 [CLS] may be favourable to using Virchow2 in practice due to the reduced storage cost for smaller embeddings; however, further task-specific validation is necessary to confirm that performance is retained in this configuration.
@@
CTransPath 0.897 0.892 0.894 0.894
Phikon 0.932 0.926 0.901 0.920
UNI 0.943 0.950 0.936 0.943
Prov-GigaPath 0.940 0.947 0.944 0.944
H-optimus-0 0.952 0.947 0.932 0.944
Virchow [CLS] 0.950 0.949 0.924 0.941
Virchow2 [CLS] 0.964 0.967 0.967 0.966
Virchow2G [CLS] 0.966 0.972 0.975 0.971
Virchow 0.950 0.948 0.933 0.944
Virchow2 0.964 0.966 0.967 0.966
Virchow2G 0.966 0.971 0.975 0.970
Table 2: Weighted F1 score results for the in-distribution tile-level benchmark tasks. Best results are bolded and second best are underlined. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only.
Model CRC
@@
MHIST PCam WILDS TILS DLBCL MIDOG Average
CTransPath 0.962 0.844 0.817 0.871 0.947 0.931 0.542 0.643 0.820
Phikon 0.959 0.888 0.794 0.904 0.972 0.944 0.544 0.700 0.838
UNI 0.963 0.943 0.843 0.934 0.983 0.945 0.565 0.749 0.866
Prov-GigaPath 0.967 0.968 0.822 0.941 0.976 0.941 0.573 0.782 0.871
H-optimus-0 0.963 0.943 0.848 0.937 0.987 0.949 0.582 0.781 0.874
Virchow [CLS] 0.970 0.932 0.836 0.934 0.966 0.948 0.591 0.760 0.867
Virchow2 [CLS] 0.976 0.971 0.860 0.935 0.985 0.947 0.615 0.800 0.886
Virchow2G [CLS] 0.974 0.970 0.857 0.944 0.983 0.947 0.628 0.804 0.888
Virchow 0.973 0.968 0.836 0.933 0.971 0.949 0.602 0.787 0.877
Virchow2 0.974 0.969 0.859 0.935 0.987 0.948 0.606 0.804 0.885
Virchow2G 0.973 0.970 0.864 0.947 0.988 0.948 0.629 0.836 0.894
Table 3: Weighted F1 score results for the out-of-distribution tile-level benchmark tasks. Best results are bolded and second best results are underlined. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only.
IDC PRAD PAAD SKCM COAD READ ccRCC HCC LUAD LIDC Avg.
CTransPath 0.476 0.363 0.416 0.557 0.118 0.093 0.177 0.063 0.560 0.255 0.308
@@
UNI 0.498 0.371 0.437 0.637 0.142 0.160 0.178 0.052 0.565 0.261 0.330
Virchow [CLS] 0.529 0.364 0.431 0.612 0.150 0.120 0.205 0.065 0.584 0.274 0.334
Virchow2 [CLS] 0.563 0.379 0.402 0.631 0.143 0.169 0.224 0.060 0.591 0.275 0.344
Virchow2G [CLS] 0.547 0.375 0.420 0.638 0.140 0.146 0.213 0.056 0.590 0.271 0.339
Virchow2G 0.559 0.385 0.458 0.632 0.139 0.175 0.222 0.062 0.588 0.274 0.350
Table 4: Pearson correlation results for the HEST-Benchmark tasks. Best results are bolded and second best results are underlined. For the Virchow models, embeddings are defined as the concatentation of class and mean patch tokens. [CLS] indicates the use of class tokens only. “Avg.” denotes the average correlation across all 10 tasks.
@@ 7 Discussion
Our ablation results demonstrate that such modifications can significantly improve benchmark performance. Changing the KoLeo regularizer to a kernel density estimator (KDE) also boosted performance. This is likely because the conditions that produce instability with KoLeo are particularly common in pathology images: that is, tiles within a minibatch are much more likely to be similar to each other compared to ImageNet [21] images within a minibatch. Furthermore, the proposed extended-context translation (ECT) is more fitting for pathology data than crop-and-resize because it avoids distorting the cell morphology which is crucial for interpreting pathology images. Although decreasing random resizing during training may reduce the model’s scale invariance, we can retain it for pathology by training across multiple magnifications. Similarly, removing solarization boosted performance (especially on PCam), as also previously observed [34, 22]. Indeed, it has been shown that the effects of augmentation are often data dependent and vary across classes [4], however, with a sufficiently large and diverse training dataset, the importance of photometric data augmentations diminishes [40]. Nevertheless, suitable image transforms, like cropping, are still necessary between the two views of an input in joint embedding models like DINOv2.
@@ A.2 Hyperparameters
Table A3: Virchow2G adapted hyperparameters from DINOv2 trained on 512 NVIDIA V100 32GB GPUs. Default augmentations are not include and are left unchanged.
A.3 Scaling plots

### v3

@@ Abstract
Foundation models are rapidly being developed for computational pathology applications. However, it remains an open question which factors are most important for downstream performance with data scale and diversity, model size, and training algorithm all playing a role. In this work, we propose algorithmic modifications, tailored for pathology, and we present the result of scaling both data and model size, surpassing previous studies in both dimensions. We introduce three new models: Virchow2, a 632 million parameter vision transformer, Virchow2G, a 1.9 billion parameter vision transformer, and Virchow2G Mini, a 22 million parameter distillation of Virchow2G, each trained with 3.1 million histopathology whole slide images, with diverse tissues, originating institutions, and stains. We achieve state of the art performance on 12 tile-level tasks, as compared to the top performing competing models. Our results suggest that data diversity and domain-specific methods can outperform models that only scale in the number of parameters, but, on average, performance benefits from the combination of domain-specific methods, data scale, and model scale.
Virchow2 Model Weights: https://huggingface.co/paige-ai/Virchow2
Virchow2 Azure AI Studio: https://aka.ms/virchow2
@@ 1 Introduction
Figure 1: Virchow2 and Virchow2G adopt a pathology-adapted training method, scale up model and data size, and increase data diversity as compared to Virchow. Gains in the average weighted F1 score across 8 tile-level benchmarks using public data (see Tab. 2) are shown with respect to two scaling axes: model size (left), in terms of the number of parameters, and dataset size (right), in terms of the number of whole slide images (WSIs). Virchow2 improves over Virchow with pathology-specific modifications to the training method while scaling up the dataset from 1.5M to 3.1M WSIs, with increased diversity. Virchow2G then scales the model size from 632M to 1.9B parameters (ViT-H to ViT-G). Benchmark performance across foundation models in pathology appears to scale with model and dataset size. Distilling Virchow2G from 1.9B parameters to 22M parameters produces Virchow2G Mini (top left), an efficient model that is competitive with Virchow (632M parameters), H-optimus-0, and GigaPath (both 1.1B parameters) at a small fraction of the parameters. For all models except CTransPath, we concatenate the class token with the mean of the patch token, improving performance.
@@ 3.1 Domain-specific augmentations
Figure 2: Highlighted modifications of DINOv2 tailored for pathology. a. Training samples are tiles drawn from the various available magnifications, e.g. 40x, 20x, 10x, or 5x. b. Illustration of the difference between standard crop-and-resize augmentation (top) and the proposed extended-context translation (ECT) augmentation (bottom). ECT uses less resizing to avoid distorting the cell morphology. c. A graphic depicting the different trends of KoLeo and KDE regularizer gradient norms as a function of distance. When samples are more likely to be similar, as is the case in pathology, large gradient norms are more likely when using KoLeo. d. Examples of common photometric augmentations used in self-supervised learning. Solarization is hypothesized to be a poor source of color invariance for pathology applications. e. Results of the ablation study where each bar reports the difference between the proposal and standard DINOv2 as measured by weighted F1 score. In the majority of instances, domain-inspired changes outperform standard approaches.
@@ 5.5 Ablation results
Fig. 2e presents the results of the ablation analysis (see Tab. A5-A8 in Appendix A.3 for additional details). In a majority of tasks (5/7), using all three domain-inspired changes, results in the best performance as well as the best average performance on in-distribution tasks. Using only ECT and KDE regularization has the best performance of 3/7 tasks and the highest average performance on out-of-distribution tasks, although the gap with all three is only 0.002. Interestingly, we observe that using only ECT leads to a small drop in performance while using only KDE regularization leads to a small improvement in performance on out-of-distribution tasks. Combining the two leads to a pronounced improvement, as noted above. All changes improved in-distribution tasks.
@@ 6 Results of large-scale training
We improve on the training recipe of Virchow with pathology specific tuning and scale up the training first along the data axis (Virchow2) and then along the model axis (Virchow2G). Afterwards, we explore the effectiveness of scaling up in order to scale down by using Virchow2G as a teacher to train a significantly smaller yet high-quality student (Virchow2G Mini). We evaluate these models against other foundation models in pathology on tile-level benchmark tasks.
@@ 6.2 Pretraining methodology
6.3 Distillation methodology
We distilled Virchow2G Mini, a ViT-S/14, from Virchow2G following the DINOv2 distillation methodology [45]. The ViT-S/14 is the standard variation of the architecture but with 4 registers and without the other modifications used in Virchow2G in order to focus on maximizing inference efficiency (i.e. GELU MLP instead of SwiGLU, no DPN, no QKN). Virchow2G Mini is trained with StableAdamW for a base learning of $2\times 10^{-4}$ with a batch size of $2048$ over a total of 1B balanced tiles. The full configuration details can be found in appendix A4.
@@ 6.4 Embedding generation
For a 224 × 224 input tile image, we choose one of two configurations of embeddings: (1) the class token (CLS-Only) and (2) the class token concatenated with the mean across all 256 of the patch tokens (CLS+Mean). The former is the recommended configuration of H-optimus-0 [47], Prov-GigaPath [60], UNI [13], and Phikon [25]; the latter is the recommended configuration for the Virchow models. We evaluate both CLS-Only and CLS+Mean embeddings for all models except CTransPath [56], for which the mean of all tokens is used as there is no class token. The CLS+Mean configurations produces embeddings of size 2,560 (1,280 $\times$ 2) for Virchow and Virchow2, 3,584 (1,792 $\times$ 2) for Virchow2G, and 768 (384 $\times$ 2) for Virchow2G Mini.
@@ 6.5 Tile benchmarks
HEST-Benchmark is a public dataset of 74 spatial transcriptomics profiles from 48 patients grouped into 10 tasks based on organ [32]. The aim of the task is to predict the expression levels of the 50 most variable genes from 112 ×$\times 112\mu$m H&E stained image patches, each centered on a spatial transcriptomics spot. Here we present the results of a random forest classifier following the protocol in [40], as opposed to the linear protocol described in Sec. 5.3.
@@ 6.6 Scaling results
We use the concatenation of the class [CLS] token with the mean of the patch tokens (CLS+Mean) as the embeddings for all Virchow models. While using the class token only (CLS) is a suggested as the default for H-optimus-0, Prov-GigaPath, UNI, and Phikon, we find that CLS+Mean embeddings produce the same or better benchmark results, especially for mitosis detection (MIDOG), a localized detection task (see Appendix A.5 for additional experimental results). Results for all tasks are shown for both embeddings configurations in Tab. 1, Tab. 2, Tab. 3, and Fig. A1 (Appendix A.4). Interestingly, Virchow2 performance is similar when using CLS or CLS+Mean embeddings, suggesting that CLS embeddings may be used with to the reduce storage cost; however, further task-specific validation is necessary to confirm that performance is retained in this configuration.
Our results show that benchmark performance scales with the size of the pre-trained model across diverse benchmarks (see Fig. A1 in Appendix A.4). Although large models are computationally expensive and their large embeddings are storage-intensive, they also offer a good starting point for model distillation. Previous work has shown DINOv2 distillation is parameter-efficient [45]. Indeed, using the same method to distill Virchow2G from a 1.9B parameter ViT-G to a 22M parameter ViT-S (roughly 1% of the parameters) we obtain a model (Virchow2G Mini) that outperforms H-optimus-0 (1.1B parameters), GigaPath (1.1B parameters), and Virchow (632M parameters) in-distribution and nearly ties Virchow out-of-distribution while outperforming H-optimus-0 (Tab. 1, Tab. 2).
Model PanMSK 20x PanMSK 10x PanMSK 5x Average
CLS+Mean Virchow2G 96.6 97.1 97.5 97.0
Virchow2 96.4 96.6 96.7 96.6
Virchow2G Mini 95.0 95.7 96.0 95.6
Virchow 95.0 94.8 93.3 94.4
H-optimus-0 95.7 95.3 93.8 95.0
Prov-GigaPath 94.2 94.8 94.8 94.6
UNI 94.4 94.8 93.8 94.3
Phikon 92.4 92.6 90.8 91.9
[…]
CLS-Only Virchow2G 96.6 97.2 97.5 97.1
Virchow2 96.4 96.7 96.8 96.6
Virchow2G Mini 95.1 96.1 96.3 95.8
Virchow 95.0 94.9 92.4 94.1
H-optimus-0 95.2 94.7 93.4 94.4
Prov-GigaPath 94.0 94.7 94.4 94.4
UNI 94.3 95.0 93.6 94.3
Phikon 92.3 92.6 90.1 91.7
NatImg 88.3 90.0 91.8 90.0
[…]
Model PCam CRC
@@
WILDS TILS MHIST DLBCL MIDOG Average
CLS+Mean Virchow2G 94.7 97.3 97.0 98.8 94.8 86.4 62.9 83.6 89.4
Virchow2 93.5 97.4 96.9 98.7 94.8 85.9 60.6 80.4 88.5
Virchow2G Mini 93.5 97.0 96.0 98.4 94.1 83.2 58.9 79.6 87.6
Virchow 93.3 97.3 96.8 97.1 94.9 83.6 60.2 78.7 87.7
H-optimus-0 93.7 96.9 92.6 98.6 94.9 84.6 59.1 79.6 87.5
Prov-GigaPath 94.1 96.9 96.9 97.9 94.2 82.3 60.3 79.4 87.8
UNI 93.5 96.7 95.4 98.3 94.6 84.8 58.4 75.9 87.2
Phikon 91.5 96.3 88.7 97.0 94.7 80.9 55.6 73.0 84.7
[…]
CLS-Only Virchow2G 94.4 97.4 97.0 98.3 94.7 85.2 62.4 80.5 88.7
Virchow2 93.5 97.6 97.1 98.5 94.7 86.0 61.5 80.0 88.6
Virchow2G Mini 93.4 97.0 96.2 98.2 93.9 82.1 59.9 76.2 87.1
Virchow 93.4 97.0 93.2 96.6 94.8 83.6 59.1 76.0 86.7
H-optimus-0 93.7 96.3 94.3 98.6 94.9 84.8 58.2 78.1 87.4
Prov-GigaPath 94.1 96.7 96.8 97.6 94.1 82.2 57.3 78.2 87.1
UNI 93.4 96.3 94.3 98.3 94.5 84.5 56.5 74.8 86.6
Phikon 90.5 95.9 88.8 97.2 94.4 79.6 54.4 69.9 83.8
NatImg 88.7 95.2 92.8 93.4 93.0 83.2 53.9 68.9 83.6
[…]
CLS+Mean Model IDC PRAD PAAD SKCM COAD READ CCRCC HCC LUAD LIDC Avg.
Virchow2G 0.559 0.385 0.458 0.632 0.139 0.175 0.222 0.062 0.588 0.274 0.349
H-optimus-0 0.555 0.388 0.469 0.650 0.158 0.161 0.222 0.062 0.596 0.278 0.354
Prov-GigaPath 0.527 0.392 0.453 0.623 0.147 0.136 0.189 0.051 0.574 0.267 0.336
UNI 0.531 0.391 0.444 0.630 0.149 0.131 0.183 0.046 0.571 0.265 0.334
Phikon 0.494 0.388 0.433 0.595 0.151 0.123 0.178 0.062 0.577 0.259 0.326
[…]
CLS-Only Virchow2G 0.547 0.375 0.420 0.638 0.140 0.146 0.213 0.056 0.590 0.271 0.339
Virchow2 0.563 0.379 0.402 0.631 0.143 0.169 0.224 0.060 0.591 0.275 0.344
Virchow 0.529 0.364 0.431 0.612 0.150 0.120 0.205 0.065 0.584 0.274 0.334
UNI 0.498 0.371 0.437 0.637 0.142 0.160 0.178 0.052 0.565 0.261 0.330
@@
[…]
@@ 7 Discussion
Our ablation results demonstrate that such modifications can significantly improve benchmark performance. Changing the KoLeo regularizer to a kernel density estimator (KDE) also boosted performance. This is likely because the conditions that produce instability with KoLeo are particularly common in pathology images: that is, tiles within a minibatch are much more likely to be similar to each other compared to ImageNet [21] images within a minibatch. Furthermore, the proposed extended-context translation (ECT) is more fitting for pathology data than crop-and-resize because it avoids distorting the cell morphology which is crucial for interpreting pathology images. Although decreasing random resizing during training may reduce the model’s scale invariance, we can retain it for pathology by training across multiple magnifications. Similarly, removing solarization boosted performance (especially on PCam), as also previously observed [35, 22]. Indeed, it has been shown that the effects of augmentation are often data dependent and vary across classes [4], however, with a sufficiently large and diverse training dataset, the importance of photometric data augmentations diminishes [42]. Nevertheless, suitable image transforms, like cropping, are still necessary between the two views of an input in joint embedding models like DINOv2. Finally, we find that different tasks benefit from different embedding configurations. High level understanding of a tissue tile can be achieved with a single “class” token that summarizes the tile; however, we find that some tasks (such as detecting cellular mitosis events) require localized information that may not be complete in this class representation. We observe that using the patch tokens in addition to the class token lead to significant performance boosts.
@@ A.2 Hyperparameters
Table A3: Virchow2G adapted hyperparameters from DINOv2 trained on 512 NVIDIA V100 32GB GPUs. Default augmentations are not included and are left unchanged.
Virchow2G Mini hyperparameter Value
Vision transformer
Patch size 14
Embedding dimension 384
Layers 12
Heads 6
MLP ratio 4
MLP activation GELU
MLP bias True
QKV bias True
Registers 4
QK normalization False
Dual PatchNorm False
Projection heads
Shared heads False
Layers 3
Bottleneck dimension 384
Hidden dimension 2048
Loss functions
Sinkhorn centering True
Student temperature 0.1
Teacher temperature 0.04
Prototypes 131072
Regularizer KDE
Regularizer parameter 5
Regularizer weight 0.05
Augmentations
Tile context size 392
Global view size 224
Local view size 98
Global crops 2
Local crops 8
Method ECT
Aspect ratio range (0.95, 1.05)
Scale range (0.9, 1.1)
Solarization False
Vertical flips True
Optimizer StableAdamW
Optimizer momentum (0.9, 0.98)
Optimizer epsilon $1\times 10^{-6}$
Optimizer weight decay (0.04, 0.2)
Learning rate $2\times 10^{-4}$
Learning rate schedule Cosine
Learning rate scaling Square Root-1024
Student EMA copy momentum 0.994
Batch size 2048
Gradient clipping norm 3.0
Drop rate 0.0
Precision FP16
Table A4: Virchow2G Mini adapted hyperparameters from DINOv2 trained on 256 NVIDIA V100 32GB GPUs. Default augmentations are not included and are left unchanged.
A.3 Detailed ablation results
Here we present the detailed results of the ablation study. Tables A5 and A6, presented the weighted F1 score for the in-domain and out-of-domain tile tasks respectively. Note all metrics have been multiplied by 100 for improved readability. The bolded numbers indicate the best performing model; in cases where multiple results are bolded, the results are not statistically significantly different at $\alpha=0.05$ as measured by McNemar’s test. We do not apply statistical testing to the average performance results and therefore do not bold any values. Tables A7 and A8 present the same analysis using only the [CLS] token as the embedding. The [CLS] token results have similar trends to those described in the main text. In general, the experiments with fewer modifications have a boost in performance when using the concatenated embedding but that delta decreases in the case where +ECT, +KDE or all three proposals are applied.
PanMSK
20x
PanMSK
10x
PanMSK
5x
ID
Average
Standard DINOv2 86.0 88.8 89.0 87.9
   +ECT 86.5 $\uparrow$ 0.5 89.4 $\uparrow$ 0.6 89.9 $\uparrow$ 0.9 88.6 $\uparrow$ 0.7
   +KDE 88.0 $\uparrow$ 2.0 90.1 $\uparrow$ 1.3 90.0 $\uparrow$ 1.1 89.4 $\uparrow$ 1.5
   -SOL 87.1 $\uparrow$ 1.1 89.2 $\uparrow$ 0.3 89.0 – 89.4 $\uparrow$ 1.5
   +ECT,+KDE 89.6 $\uparrow$ 3.6 92.1 $\uparrow$ 3.2 92.2 $\uparrow$ 3.2 91.3 $\uparrow$ 3.4
   +ECT,+KDE,-SOL 89.9 $\uparrow$ 3.9 92.4 $\uparrow$ 3.6 93.3 $\uparrow$ 4.3 91.9 $\uparrow$ 4.0
Table A5: Weighted F1 score (multiplied by 100) for the in-domain tasks. ECT is the proposed extended context translation, KDE is the proposed regularization function and SOL refers to the removal of solarization.
PCam CRC TILS MHIST MIDOG
OOD
Average
Standard DINOv2 83.6 94.0 92.5 79.4 63.0 82.5
   +ECT 84.9 $\uparrow$ 1.3 93.3 $\downarrow$ 0.7 92.6 $\uparrow$ 0.1 76.2 $\downarrow$ 3.2 63.6 $\uparrow$ 0.6 82.1 $\downarrow$ 0.5
   +KDE 84.1 $\uparrow$ 0.6 94.9 $\uparrow$ 0.9 93.3 $\uparrow$ 0.8 78.3 $\downarrow$ 1.0 66.5 $\uparrow$ 3.5 83.4 $\uparrow$ 1.0
   -SOL 84.8 $\uparrow$ 1.3 93.8 $\downarrow$ 0.2 92.7 $\uparrow$ 0.2 77.2 $\downarrow$ 2.2 65.9 $\uparrow$ 2.9 82.9 $\uparrow$ 0.4
   +ECT, +KDE 86.7 $\uparrow$ 3.1 95.3 $\uparrow$ 1.3 93.0 $\uparrow$ 0.5 80.3 $\uparrow$ 0.9 66.5 $\uparrow$ 3.6 84.4 $\uparrow$ 1.9
   +ECT, +KDE, -SOL 86.7 $\uparrow$ 3.2 95.8 $\uparrow$ 1.8 93.2 $\uparrow$ 0.7 79.2 $\downarrow$ 0.2 66.3 $\uparrow$ 3.4 84.3 $\uparrow$ 1.8
Table A6: Weighted F1 score (multiplied by 100) for the out-of-domain tasks. ECT is the proposed extended context translation, KDE is the proposed regularization function and SOL refers to the removal of solarization.
PanMSK
20x
PanMSK
10x
PanMSK
5x
ID
Average
Standard DINOv2 83.2 86.0 86.3 85.1
   +ECT 82.7 $\downarrow$ 0.4 86.1 $\uparrow$ 0.1 85.9 $\downarrow$ 0.3 84.9 $\downarrow$ 0.2
   +KDE 87.4 $\uparrow$ 4.2 89.8 $\uparrow$ 3.8 89.6 $\uparrow$ 3.3 88.9 $\uparrow$ 3.8
   -SOL 84.4 $\uparrow$ 1.2 87.0 $\uparrow$ 1.0 86.9 $\uparrow$ 0.6 86.1 $\uparrow$ 1.0
   +ECT,+KDE 89.3 $\uparrow$ 6.2 91.9 $\uparrow$ 5.8 92.1 $\uparrow$ 5.8 91.1 $\uparrow$ 6.0
   +ECT, +KDE, -SOL 89.7 $\uparrow$ 6.6 92.4 $\uparrow$ 6.4 93.1 $\uparrow$ 6.9 91.8 $\uparrow$ 6.7
Table A7: Weighted F1 score (multiplied by 100) for the in-domain tasks where the embedding is [CLS] only. ECT is the proposed extended context translation, KDE is the proposed regularization function and SOL refers to the removal of solarization.
PCam CRC TILS MHIST MIDOG
OOD
Average
Standard DINOv2 83.2 92.1 91.7 80.1 62.1 81.8
   +ECT 82.3 $\downarrow$ 0.9 91.3 $\downarrow$ 0.8 91.1 $\downarrow$ 0.6 74.0 $\downarrow$ 6.1 63.3 $\uparrow$ 1.2 80.4 $\downarrow$ 1.4
   +KDE 84.2 $\uparrow$ 1.0 95.8 $\uparrow$ 3.7 93.1 $\uparrow$ 1.4 81.3 $\uparrow$ 1.3 66.4 $\uparrow$ 4.2 84.2 $\uparrow$ 2.4
   -SOL 83.3 $\uparrow$ 0.1 92.9 $\uparrow$ 0.8 92.2 $\uparrow$ 0.5 77.5 $\downarrow$ 2.6 65.5 $\uparrow$ 3.3 82.3 $\uparrow$ 0.5
   +ECT,+KDE 86.6 $\uparrow$ 3.4 96.3 $\uparrow$ 4.3 92.9 $\uparrow$ 1.2 79.9 $\downarrow$ 0.2 67.4 $\uparrow$ 5.2 84.6 $\uparrow$ 2.8
   +ECT,+KDE,-SOL 87.3 $\uparrow$ 4.2 96.9 $\uparrow$ 4.9 93.1 $\uparrow$ 1.4 78.8 $\downarrow$ 1.3 65.6 $\uparrow$ 3.5 84.4 $\uparrow$ 2.6
Table A8: Weighted F1 score (multiplied by 100) for the out-of-domain tasks where the embedding is [CLS] only. ECT is the proposed extended context translation, KDE is the proposed regularization function and SOL refers to the removal of solarization.
A.4 Scaling plots